The half-life of cobalt-60 is 5.274 years. Calculate the activity of a sample of
cobalt-60 weighing 1 g in units of Curie (Ci).
1
Expert's answer
2020-03-27T10:51:32-0400

"N=N_oe^{-\lambda t}"

activity is "\frac{dN}{dt}"=A


"A_t=A_oe^{-\lambda t}"

"\lambda=\frac{\ln2}{t_{1/2}}=\frac{0.69}{5.274\times365\times24\times60\times60}=\frac{.69}{0.166\times10^9}=4.16\times10^{-9}s^{-1}"

"A_o=\lambda N_o=4.16\times10^{-9}\times\frac{1}{60}\times6.023\times10^{23}" disintegration/sec

"A_o=4.17\times10^{14}" disintegration/sec

1 Ci=3.7x1010 disintegration/sec

converting Ao into curie

Ao = 1.13x104 Ci